# Deep Learning Based Dynamics Identification and Linearization of Orbital Problems using Koopman Theory

### v2

@@ Abstract
The study of the Two-Body and Circular Restricted Three-Body Problems in the field of aerospace engineering and sciences is deeply important because they help describe the motion of both celestial and artificial satellites. With the growing demand for satellites and satellite formation flying, fast and efficient control of these systems is becoming ever more important. Global linearization of these systems allows engineers to employ methods of control in order to achieve these desired results. We propose a data-driven framework for simultaneous system identification and global linearization of both the Two-Body Problem and Circular Restricted Three-Body Problem via deep learning-based Koopman Theory, i.e., a framework that can identify the underlying dynamics and globally linearize it into a linear time-invariant (LTI) system. The linear Koopman operator is discovered through purely data-driven training of a Deep Neural Network with a custom architecture. This paper displays the ability of the Koopman operator to generalize to various other Two-Body systems without the need for retraining. We also demonstrate the capability of the same architecture to be utilized to accurately learn a Koopman operator that approximates the Circular Restricted Three-Body Problem.
@@ 1 INTRODUCTION
The rapid growth in motivation to explore space in the last fifteen years has sparked a number of significant missions to planets such as Mars and Venus, moons such as Enceladus and Titan, as well as our own Moon. From interplanetary missions to the placement of satellites within our own planet’s orbit, we would not be able to achieve the goals of these endeavors without understanding and exploiting the natural dynamics of these systems. Much effort has been spent in determining how exactly the dynamics of bodies, either natural or man-made, are influenced, and how engineers can develop guidance, navigation, and control systems to complete these extraordinary tasks. Today, thousands of satellites orbit around our Earth in a variety of different orbits, and the governing equations of motion for all of these satellites are the equations originating from the N-Body Problem. In the N-Body equations of motion, the only force that imparts energy onto the bodies is gravity [1], and dependent on the complexity of the solution required and the nature of the system in question, the number of bodies varies. Most often in celestial dynamics, the Two-Body Problem (2BP) and Circular Restricted Three-Body Problem (CR3BP) are analysed and used [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. As we endeavour to reach the Moon again, and move forth to Mars, the analysis and use of the CR3BP will become even more prevalent [7, 10, 11, 12, 13].
@@
The case of the CR3BP is a well-known and studied reduction of the Three-Body problem, to which no global analytical solutions exist [7, 10, 11]. Instead, quasi-periodic, periodic and hyperbolic invariant manifolds present themselves as special solutions, which can inform us on the dynamics of the CR3BP [5, 17, 7, 10, 11]. This system is of importance because the Earth-Moon system can be treated as such [8], with the orbit of the Moon around the Earth having an eccentricity of 0.0549 [18]. The Earth-Moon system has five equilibrium points, the so-called Lagrange points [9], around which it is possible to generate periodic orbits that are intergrable. Much like the 2BP, efficient, accurate linear representation of the CR3BP would further advance the exploration of space, by reducing the complexity of creating control systems capable of being used in these systems.
@@
Koopman theory, first proposed in 1931 [19], has gained traction over the last few years as a solution for finding global linearization of nonlinear systems. In a nutshell, the theory states that the dynamics of a nonlinear system can be described linearly by an infinite-dimensional Koopman operator. Due to its infinite dimension, for practical use, it is typically approximated using data-driven methods such as Extended Dynamic Mode Decomposition (EDMD) [20]. The process of implementing EDMD on the nonlinear system to find an accurate approximation is not only highly involved, but requires an innate understanding of the method in order to select the basis functions. These basis functions are the building blocks from which functions known as observable functions are constructed. These observable functions are then what is applied to the data in order to lift dynamics to the appropriate dimension. However, constructing a feed-forward neural network (NN) that implements EDMD makes it possible to find an approximate general Koopman operator linearization applicable to a region of state space, that is time invariant and constant. Because of the high interest in Koopman Theory as a method for globally linearizing the nonlinear dynamics of a problem, a great deal of research has been conducted in a wide variety of fields. Applications in fluid dynamics [21], robotics [22] and control systems [23] have all been investigated. Whilst the utilization of Koopman in order to develop linear systems capable of control has been researched heavily, with many models including optimal controls [24], Model Predictive Controllers (MPC) [25], Linear Quadratic Controllers (LQR) [26], all being successfully developed through the implementation of EDMD with control (EDMDc) [27] via analytical solutions or the development of NN’s. Another advantage of using a data-driven method is that it does not require system knowledge, and the nonlinear system can be completely unknown [28, 29].
@@
Neural Networks, although having a history that spans almost 40 years, have found a resurgence in the last 10-15 years, propelled by the work of LeCun et al. [30], who were able to analyse the ImageNet dataset and accurately categorize image classes within the dataset. Since then, the advancements of big data coupled with computational power has allowed Deep Neural Networks (DNNs) to grow rapidly, and be integrated into many fields of research and development. The power of DNN’s lies in their ability to represent any arbitrary function, including the Koopman observable functions required to linearize the nonlinear dynamics [31]. This is possible through a DNN with an adequate number of hidden neurons and a linear output layer, satisfying the universal approximation theorem [32, 33, 34]. The use of a DNN in the approximation of the Koopman operator relieves us from the need to determine a satisfactory set of basis functions to which the Koopman eigenfucntions are mapped from, which is often difficult and can lead to poor approximations if not chosen correctly.
@@
Current state of the art of the linearization of the TBP and CR3BP includes the use of Koopman Operator Theory. The relative, Clohessy-Wiltshire equations were successfully modelled by Koopman Theory to relieve some of the underlying assumptions as shown by [37]. Other models, including an oblate planet problem [38], Sun-synchronous orbit [39] and Earth-Moon L1 Lagrange point system [40] have all been successfully linearized with the Koopman operator. On the other hand, [41, 2, 42] were able to utilize the Koopman operator to study and anaylze the separate challenges related to modelling the CR3BP, including models around colinear Lagrange points and uncertainties in the propagation of the dynamics. The work [43] however, implements EDMD to analyse the CR3BP in regions outside of these Lagrange points, expanding on the breadth of knowledge base of Koopman applications on orbital problems. Because of the numerical complexity of both the TBP and CR3BP, the treatment of the theory in this domain is often contextualised to specific conditions rather than generalized to the entire problem as a whole.
@@
This work extends the framework developed in Tiwari et al. [44]. The framework of the DNN developed by Tiwari et al. is kept the same, with improvements added for this application. Although the application included control, it is not incorporated for the purposes of this paper when the methodology is applied to both the 2BP and CR3BP. To the best of our knowledge, this is the first time that both the 2BP and the CR3BP have been globally linearized with reasonable precision.
@@
    First we demonstrate the ability of the learned Koopman operator to globally linearize a Circular Two-Body problem centered around the Earth for orbits with an apogee altitude ranging between 200-30000km. We also demonstrate the ability of the same network, without additional training, to generalize to other two body systems, specifically around the Moon and Jupiter. We demonstrate that the approximation is accurate and that various invariant properties of the circular orbit are conserved.
@@
    Second, we demonstrate the applicability of the Koopman operator to the more complicated Circular-Restricted Three-Body Problem (CR3BP). We globally linearise the periodic orbit of a satellite that orbits the Earth, including the influence of the Moon, with the initial position of the satellite being close to the L1 Lagrange point. We also analyze the accuracy of the approximation by studying the conservation of the Jacobi constant in the CR3BP, showing that our model can adequately capture the value of this constant throughout the evolution of the dynamics.
@@ 2.1.1 Koopman Operator
where, $\boldsymbol{x}(t)\in\mathbb{R}^{n}$ is the system state at time $t$ and $\boldsymbol{F}$ is the function that describes the evolution of the state in the continuous sense. This continuous system can also be modelled in a discrete-time representation by evaluating the solution to the system at finite, discrete time intervals $\Delta t$, such that $\boldsymbol{x}_{k}=x(k\Delta t).$ Hence, the discrete-time dynamics are represented as:
@@ 2.1.2 Dynamic Mode Decomposition and its Extensions
Determining the set of observable functions that span a Koopman invariant subspace, required to lift the states, is not a trivial task. Hence, a number of methods have been derived such as Dynamic Mode Decomposition (DMD), which aims to extract the dynamic modes of a given set of data. These modes can be interpreted as the generalization of the global stability modes and therefore project the underlying physcial mechanisms of the system [46].
@@ 2.1.3 Koopman Approximation Algorithm
In this work, the finite dimension Koopman operator is approximated using the combination of EDMD with deep neural networks (DNNs). The procedure in which this is achieved is based off of our method in our previous work, [44], with the work presented herein being an extension of the architecture and structure presented previously. The DNN is used in order to learn and define the set of observables that are used to lift the states, contrary to other methods in which these observable functions are selected manually as a dictionary [48, 49, 50]. Subsequently, the finite-dimensional approximation to the Koopman operator is calculated by utilizing least-squares regression (see Algorithm 1) [44].
@@
Because the Koopman operator calculated in this approach represents an approximation to the real, infinite dimensional operator, and as the observable functions do not span a Koopman invariant subspace, the predicted state is denoted with a circumflex (hat) symbol to emphasise that it is an approximation:
@@
2.2 Circular Orbital Dynamics
One of the two main applications of the proposed model in this work is to linearize a circular 2BP orbit. Orbiting satellites in Low Earth Orbit (LEO) range in altitude from 200-300km to around 2000km [55], and have an eccentricity of less than 0.25, whilst Geostationary orbiting satellites have an ideal eccentricity of 0 [56]. With many of the operational satellites having an eccentricity so close to zero, this aspect motivates us to investigate the linearization of those orbits that are circular, i.e., with zero eccentricity. The equations of motion for a satellite in any orbit is a second-order differential equation that relates the movement of the satellite to the accelerations that perturb the orbit. These accelerations can depend on a number of parameters such as time $t$, position $\vec{\mathbf{r}}$, velocity $\vec{\mathbf{v}}$ or physical forces $\vec{\mathbf{f}}$.
@@ 2.2 Circular Orbital Dynamics
Where, $G$ is the Gravitational Constant, $M$ is the mass of the central body, $m$ is the mass of the orbiting body (satellite), $\vec{\mathbf{R}}_{2}-\vec{\mathbf{R}}_{1}$ is the distance from center of the satellite to the central body and $\mathbf{a}$ is the additional perturbing acceleration. In our work, we set $\mathbf{a}=\vec{\mathbf{0}}$ to preserve the circularity of the orbit, which should improve the DNN’s ability to learn the observables, consequently leading to a more accurate representation of the Koopman operator. To prove the ability of the model, we simplify the nonlinear dynamics by choosing not to include orbital perturbations such as drag, solar radiation pressure, or a complex gravity model such as the J2 perturbation model, but they can be included in the nonlinear dynamics, at the expense of a potentially larger K matrix. We also set the position of the central body, $\vec{\mathbf{R}}_{1}$, to be at the origin, hence simplifying 18 to:
@@
To ease the generation of data for this system, each orbit obtains the same initial pose, being,  $x=r,y=0$. Hence, the initial velocity is purely tangential, resulting in the initial velocities being:
@@
$\dot{y}=\sqrt{\frac{\mu}{r}}$
@@
$\mathbf{x_{0}}=\begin{bmatrix}r&0&0&\sqrt{\frac{\mu}{r}}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$ (23)
Where, $r$ is the radius of the given circular orbit, which is a randomly selected parameter for each initial condition in the training dataset.
@@ 2.3 Circular Restricted Three-Body Dynamics
In order to produce trajectories that can be used in training with enough variance to adequately train the DNN, the initial position in the $x$ direction was multiplied by a random number in the range $[1,1.05]$. Such a small range is required since the nature of the orbit and its periodic behaviour are highly sensitive to the initial conditions, a manifestation of chaotic behavior. Multipliers of any value outside of this narrow range would result in orbits that were not periodic around the L1 point.

### v3

@@ Abstract
The study of the Two-Body and Circular Restricted Three-Body Problems in the field of aerospace engineering and sciences is deeply important because they help describe the motion of both celestial and artificial satellites. With the growing demand for satellites and satellite formation flying, fast and efficient control of these systems is becoming ever more important. Global linearization of these systems allows engineers to employ methods of control in order to achieve these desired results. We propose a data-driven framework for simultaneous system identification and global linearization of the Circular, Elliptical and Perturbed Two-Body Problem as well as the Circular Restricted Three-Body Problem around the L1 Lagrange point via deep learning-based Koopman Theory, i.e., a framework that can identify the underlying dynamics and globally linearize it into a linear time-invariant (LTI) system. The linear Koopman operator is discovered through purely data-driven training of a Deep Neural Network with a custom architecture. This paper displays the ability of the Koopman operator to generalize to various other Two-Body systems without the need for retraining. We also demonstrate the capability of the same architecture to be utilized to accurately learn a Koopman operator that approximates the Circular Restricted Three-Body Problem.
@@ 1 INTRODUCTION
The rapid growth in motivation to explore space in the last fifteen years has sparked a number of significant missions to planets such as Mars and Venus, moons such as Enceladus and Titan, as well as our own Moon. From interplanetary missions to the placement of satellites within our own planet’s orbit, we would not be able to achieve the goals of these endeavors without understanding and exploiting the natural dynamics of these systems. Much effort has been spent in determining how exactly the dynamics of bodies, either natural or man-made, are influenced, and how engineers can develop guidance, navigation, and control systems to complete these extraordinary tasks. Today, thousands of satellites orbit around our Earth in a variety of different orbits, and the governing equations of motion for all of these satellites are the equations originating from the N-Body Problem. In the N-Body equations of motion, the only force that imparts energy onto the bodies is gravity [1], and dependent on the complexity of the solution required and the nature of the system in question, the number of bodies varies. Most often in celestial dynamics, the Two-Body Problem (2BP) and Circular Restricted Three-Body Problem (CR3BP) are analyzed and used [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. As we endeavor to reach the Moon again, and move forth to Mars, the analysis and use of the CR3BP will become even more prevalent [7, 10, 11, 12, 13].
@@
The case of the CR3BP is a well-known and studied reduction of the Three-Body problem, to which no global analytical solutions exist [7, 10, 11]. Instead, quasi-periodic, periodic and hyperbolic invariant manifolds present themselves as special solutions, which can inform us on the dynamics of the CR3BP [5, 17, 7, 10, 11]. This system is of importance because the Earth-Moon system can be treated as such [8], with the orbit of the Moon around the Earth having an eccentricity of 0.0549 [18]. The Earth-Moon system has five equilibrium points, the so-called Lagrange points [9], around which it is possible to generate periodic orbits that are integrable. Much like the 2BP, efficient, accurate linear representation of the CR3BP would further advance the exploration of space, by reducing the complexity of creating control systems capable of being used in these systems.
@@
Koopman theory, first proposed in 1931 [19], has gained traction over the last few years as a solution for finding global linearization of nonlinear systems. In a nutshell, the theory states that the dynamics of a nonlinear system can be described linearly by an infinite-dimensional Koopman operator. Due to its infinite dimension, for practical use, it is typically approximated using data-driven methods such as Extended Dynamic Mode Decomposition (EDMD) [20]. The process of analytically implementing these approximating methods such as EDMD or Hankel-DMD [21] on the nonlinear system to find an accurate approximation is not only highly involved, but requires an innate understanding of the method in order to select a reasonable, diverse and encompassing dictionary of basis functions. These basis functions are the building blocks from which functions known as observable functions are constructed. These observable functions are then what is applied to the data in order to lift dynamics to the appropriate dimension. However, the main focus of our work is in the construction of a feed-forward neural network (NN) that implements EDMD, making it easier to find an approximate general Koopman operator linearization applicable to a region of state space, that is time invariant and constant. Because of the high interest in Koopman Theory as a method for globally linearizing the nonlinear dynamics of a problem, a great deal of research has been conducted in a wide variety of fields. Applications in fluid dynamics [22], robotics [23] and control systems [24] have all been investigated. Whilst the utilization of Koopman in order to develop linear systems capable of control has been researched heavily, with many models including optimal controls [25], Model Predictive Controllers (MPC) [26], Linear Quadratic Controllers (LQR) [27], all being successfully developed through the implementation of EDMD with control (EDMDc) [28] via analytical solutions or the development of NN’s. Another advantage of using a data-driven method is that it does not require system knowledge, and the nonlinear system can be completely unknown [29, 30].
@@
Neural Networks, although having a history that spans almost 40 years, have found a resurgence in the last 10-15 years, propelled by the work of LeCun et al. [31], who were able to analyze the ImageNet dataset and accurately categorize image classes within the dataset. Since then, the advancements of big data coupled with computational power has allowed Deep Neural Networks (DNNs) to grow rapidly, and be integrated into many fields of research and development. The power of DNN’s lies in their ability to represent any arbitrary function, including the Koopman observable functions required to linearize the nonlinear dynamics [32]. This is possible through a DNN with an adequate number of hidden neurons and a linear output layer, satisfying the universal approximation theorem [33, 34, 35]. The use of a DNN in the approximation of the Koopman operator relieves us from the need to determine a satisfactory set of basis functions to which the Koopman eigenfucntions are mapped from, which is often difficult and can lead to poor approximations if not chosen correctly.
@@
Current state of the art of the linearization of the 2BP and CR3BP includes the use of Koopman Operator Theory. The relative, Clohessy-Wiltshire equations were successfully modeled by Koopman Theory to relieve some of the underlying assumptions as shown by [38]. Other models, including an oblate planet problem [39], Sun-synchronous orbit [40] and Earth-Moon L1 Lagrange point system [41] have all been successfully linearized with the Koopman operator. On the other hand, [42, 2, 43] were able to utilize the Koopman operator to study and analyze the separate challenges related to modeling the CR3BP, including models around co-linear Lagrange points and uncertainties in the propagation of the dynamics. The work [44] however, implements EDMD to analyze the CR3BP in regions outside of these Lagrange points, expanding on the breadth of knowledge base of Koopman applications on orbital problems. Because of the numerical complexity of both the 2BP and CR3BP, the treatment of the theory in this domain is often contextualized to specific conditions rather than generalized to the entire problem as a whole.
@@
This work extends the framework developed in Tiwari et al. [45]. The framework of the DNN developed by Tiwari et al. is kept the same, with improvements added for this application, being the modification of the loss function to be more suited to the significantly more complex and nonlinear dynamics that is the 2BP and CR3BP. We emphasize the significance of the CR3BP dynamics and the simplicity of our framework that is able to provide a global linearization around a L1 point. Although the previous application included control, it is not incorporated for the purposes of this paper when the methodology is applied to both the 2BP and CR3BP. To the best of our knowledge, this is the first time that both the 2BP and the CR3BP have been globally linearized with reasonable precision.
@@
    First we demonstrate the ability of the learned Koopman operator to globally linearize both the Circular and Elliptical Two-Body problem centered around the Earth for orbits with an apogee altitude ranging between 200-30,000 km. We also demonstrate the ability of the same network, without additional training, to generalize to other two body systems, specifically around the Moon and Jupiter. We demonstrate that the approximation is accurate and that various invariant properties of the circular orbit are conserved. We also demonstrate that the prediction is reasonably accurate for significantly extended time horizon predictions, namely for 10 orbital periods. And we also show the capability of our model to capture non-periodic perturbations such as J2 and Solar Radiation Pressure (SRP).
@@
    Second, we demonstrate the applicability of the Koopman operator to the more complicated Circular-Restricted Three-Body Problem (CR3BP). We globally linearize the periodic orbit of a satellite that orbits the Earth, including the influence of the Moon, with the initial position of the satellite being close to the L1 Lagrange point. We also analyze the accuracy of the approximation by studying the conservation of the Jacobi constant in the CR3BP, showing that our model can adequately capture the value of this constant throughout the evolution of the dynamics.
@@ 2.1.1 Koopman Operator
where, $\boldsymbol{x}(t)\in\mathbb{R}^{n}$ is the system state at time $t$ and $\boldsymbol{F}$ is the function that describes the evolution of the state in the continuous sense. This continuous system can also be modeled in a discrete-time representation by evaluating the solution to the system at finite, discrete time intervals $\Delta t$, such that $\boldsymbol{x}_{k}=x(k\Delta t).$ Hence, the discrete-time dynamics are represented as:
@@ 2.1.2 Dynamic Mode Decomposition and its Extensions
Determining the set of observable functions that span a Koopman invariant subspace, required to lift the states, is not a trivial task. Hence, a number of methods have been derived such as Dynamic Mode Decomposition (DMD), which aims to extract the dynamic modes of a given set of data. These modes can be interpreted as the generalization of the global stability modes and therefore project the underlying physical mechanisms of the system [47].
@@ 2.1.3 Koopman Approximation Algorithm
In this work, the finite dimension Koopman operator is approximated using the combination of EDMD with deep neural networks (DNNs). The procedure in which this is achieved is based off of our method in our previous work, [45], with the work presented herein being an extension of the architecture and structure presented previously with physics-informed modifications to the loss function, in order to better suit the problem. The DNN is used in order to learn and define the set of observables that are used to lift the states, contrary to other methods in which these observable functions are selected manually as a dictionary [49, 50, 51]. This enables us to save significant time by not requiring the determination of a dictionary of basis functions, which need to be hand crafted for every problem, that serve to create the observable functions. Subsequently, the finite-dimensional approximation to the Koopman operator is calculated by utilizing least-squares regression (see Algorithm 1) [45].
@@
Because the Koopman operator calculated in this approach represents an approximation to the real, infinite dimensional operator, and as the observable functions do not span a Koopman invariant subspace, the predicted state is denoted with a circumflex (hat) symbol to emphasize that it is an approximation:
@@
2.2 Orbital Dynamics
One of the two main applications of the proposed model in this work is to linearize both the circular and elliptical 2BP orbit. Orbiting satellites in Low Earth Orbit (LEO) range in altitude from 200-300 km to around 2000 km [56], and have an eccentricity of less than 0.25, whilst Geostationary orbiting satellites have an ideal eccentricity of 0 [57]. With many of the operational satellites having an eccentricity less than 0.5 or very close to zero, this aspect motivates us to investigate the linearization of those orbits that are within this eccentricity range. The equations of motion for a satellite in any orbit is a second-order differential equation that relates the movement of the satellite to the accelerations that perturb the orbit. These accelerations can depend on a number of parameters such as time $t$, position $\vec{\mathbf{r}}$, velocity $\vec{\mathbf{v}}$ or physical forces $\vec{\mathbf{f}}$.
@@ 2.2 Orbital Dynamics
Where, $G$ is the Gravitational Constant, $M$ is the mass of the central body, $m$ is the mass of the orbiting body (satellite), $\vec{\mathbf{R}}_{2}-\vec{\mathbf{R}}_{1}$ is the distance from center of the satellite to the central body and $\mathbf{a}$ is the additional perturbing acceleration. In our work, we set $\mathbf{a}=\vec{\mathbf{0}}$ when dealing with purely circular or elliptical orbits to preserve the simplicity of the orbit, which should improve the DNN’s ability to learn the observables, consequently leading to a more accurate representation of the Koopman operator. We also set the position of the central body, $\vec{\mathbf{R}}_{1}$, to be at the origin, hence simplifying 18 to:
@@
To prove the ability of the model to also handle some of these perturbing accelerations, we perform simulations which include orbital perturbations such as solar radiation pressure and a complex gravity model such as the J2 perturbation. The models for both perturbations can be found in [58] and for these simulations, the nonlinear dynamics can be represented as follows:
$\vec{\ddot{\mathbf{r}}}=\mu\frac{\vec{\mathbf{r}}}{\left|\vec{\mathbf{r}}\right|^{3}}+\vec{\mathbf{a}}_{J2}+\vec{\mathbf{a}}_{SRP}$ (20)
@@
To ease the generation of data for this system, each orbit obtains the same initial pose, chosen to be at the periapsis of the orbit, $x=r_{p},\>y=0$. Hence, the initial velocity is purely tangential, and defined as:
@@
$\dot{y}=\sqrt{\mu\left(\frac{2}{r_{p}}-\frac{1}{a}\right)}$
Where, $r_{p}$ is the radius of periapsis for the orbit and $a$ is the semi-major axis. For purely circular orbits this simplifies to: $\dot{y}=\sqrt{\frac{\mu}{r_{p}}}$.
@@
$\mathbf{x_{0}}=\begin{bmatrix}r_{p}&0&0&\sqrt{\mu\left(\frac{2}{r_{p}}-\frac{1}{a}\right)}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$ (24)
and $r_{p}$ is a randomly selected parameter for each initial condition in the training dataset.
@@ 2.3 Circular Restricted Three-Body Dynamics
In order to produce trajectories that can be used in training with enough variance to adequately train the DNN, the initial position in the $x$ direction was multiplied by a random number in the range $[1,1.05]$. Such a small range is required since the nature of the orbit and its periodic behavior are highly sensitive to the initial conditions, a manifestation of chaotic behavior. Multipliers of any value outside of this narrow range would result in orbits that were not periodic around the L1 point.